OpenStudy (anonymous):

WILL MEDAL Solve each equation. Check your answer: 4 &1/4= 1 &3/4 + p

Mehek (mehek14):

\(\tt{4\dfrac{1}{4}=1\dfrac{3}{4}+p}\) <- like that?

OpenStudy (anonymous):

yes

Mehek (mehek14):

turn both numbers into improper fractions do you know how to do that?

OpenStudy (anonymous):

cant really remember right now. eek.

Mehek (mehek14):

4*4 = 16 16 + 1 = 17 \(\dfrac{17}{4}\)

Mehek (mehek14):

multiply the whole number by the denominator and add the numerator

OpenStudy (anonymous):

o.k. just a second.

OpenStudy (anonymous):

o.k. 17/4 - 7/4 = 2.5 or 5/2. now what

Mehek (mehek14):

o.O where did you get the other numbers from?

OpenStudy (anonymous):

whattttttt?

Mehek (mehek14):

\(\tt{4\dfrac{1}{4}=\dfrac{17}{4}\\1\dfrac{3}{4}=\dfrac{7}{4}}\) <- like that?

Mehek (mehek14):

Oh wait you solved it

Mehek (mehek14):

my bad I didn't see that you are correct p = 5/2

OpenStudy (anonymous):

thanks.

Mehek (mehek14):

yw